OpenStudy (anonymous):

Graph the hyperbola with equation ((x+1)^2/9)-((y+4)^2/25) = 1

OpenStudy (anonymous):

@nincompoop

OpenStudy (aum):

What is the center of this hyperbola? What are the 'a' and 'b' values? Compare it to \(\Large \frac{(x−h)^2}{a^2}−\frac{(y−k)^2}{b^2}=1 \) where (h,k) is the center.

OpenStudy (anonymous):

a=3 b=5 and then (-1,-4)?

OpenStudy (aum):

Yes. Plot the point (-1, -4). That will be the center of the hyperbola. Draw a horizontal line from the center and on it mark 'a' units to the right and 'a' units to the left. Those will be the vertex. Then draw the hyperbola.

OpenStudy (anonymous):

 thats a reaaallly bad drawing but kind of like that?